OpenStudy (anonymous):

You and two friends visit a pizza shop. There are 13 toppings. 8 toppings are meat and the rest are vegetable. Each of you orders a different topping (no repeats) at random. What is the probability that your group orders only meat toppings?

OpenStudy (jchick):

The first person has a 8/13 chance of getting a meat topping. The second person has a 7/12 chance because there are no repeats, which means the third person would have a 6/11 chance. Since there are three of you, you multiply 8/13 * 7/12 * 6/11. That gives you a 19.6% chance or 336/1716 Odds for first person to pick a meat topping: 8/13 Odds for second person to pick meat: 7/12 Odds for third person to pick meat: 6/11 Total odds = 8/13 * 7/12 * 6/11 = 336 / 1716 = 28 / 143 = 19.58%

OpenStudy (jchick):

28/143

OpenStudy (jchick):

(8/13)*(7/13)*(6/13)